OpenStudy (anonymous):

heres a kindergarden question solve (8-1*6)+(18/9)-(6/4)=?

OpenStudy (anonymous):

Did you need help with this? P E M D A S

OpenStudy (mathmate):

Note that the multiplication is carried out before addition!

OpenStudy (anonymous):

no just for fun

OpenStudy (hoblos):

=2+ 2 - 3/2 = 5/2

OpenStudy (anonymous):

was that the answer?

OpenStudy (anonymous):

i think you over thought it